OpenStudy (eric_d):

Show that sin 3 theta = 3 sin theta - 4 sin^3 theta.

OpenStudy (eric_d):

Hence, show that sin 3 theta- cos 2 theta = (1-s)(4s^2+2s-1), where s = sin theta

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

use identety of sin(A+B) since 3=2+1

OpenStudy (eric_d):

That's all

OpenStudy (ikram002p):

simplify

OpenStudy (ikram002p):

lets do it together :)

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

sin(A+B) =sin A cos B + cos A sin B =

OpenStudy (eric_d):

I'm nt sure

OpenStudy (eric_d):

Hw do I begin

OpenStudy (ikram002p):

\(\large \sin 3 \theta=\sin (2 \theta +\theta)=\sin 2 \theta \cos \theta +\cos 2\theta\sin \theta \)

OpenStudy (eric_d):

ok what do I need to simplify

OpenStudy (ikram002p):

http://prntscr.com/4gwqnr

OpenStudy (eric_d):

 like this ?